### v1

@@ 1 Introduction
Machine learning (ML) is revolutionising different industries in recent years. Despite its widespread adoption, a key challenge in deploying ML is ensuring data privacy. One promising direction is to employ advance cryptosystems such as multiparty computation (MPC) and homomorphic encryption (HE) during model training and model inference, resulting in privacy-preserving training and privacy-preserving inference. In this paper, we use the term Privacy-preserving machine learning (PPML) to cover both cases. While in principle these advance cryptosystems can be used to protect sensitive data used in the computation of any functionality, in practice they are designed to support basic arithmetics operations (i.e., addition and multiplication) or basic Boolean operations (i.e., AND and OR) only. This limitation makes it expensive to compute non-polynomial functions such as ReLU, sigmoid, and maxpool, functions commonly found in ML, and in particular, neural networks. Existing PPML address this limitations by replacing these functions with their polynomial approximations. In this paper, we refer to this type of neural network in which non-polynomial functions have been replaced with polynomials as polynomial approximation of neural network (PANN). Similar to existing MPC and HE based PPML, this paper focuses on privacy-preserving inference since privacy-preserving training using HE and MPC are still considered too inefficient in practice.
Figure 1: Time cost for our models and the state-of-the-art to achieve the previous best accuracy on PANN (CIFAR-10).
Naturally, the (implicit) use of PANN in privacy-preserving inference involves two design considerations, namely, the precision of the polynomial approximation and the accuracy of the PANN (compared with the underlying backbone model). Intuitively, the accuracy of inference cannot be guaranteed if the approximation introduces a large error. Early schemes often adopt simple polynomials in approximation. For example, CryptoNets and HCNN utilized the square function to replace activation functions Gilad-Bachrach et al. (2016); Badawi et al. (2020). CryptoDL used degree 2 and degree 3 polynomials to approximate the ReLU function Hesamifard et al. (2017), while Faster Cryptonets leveraged minimax approximation with degree two Chou et al. (2018). nGraph-HE and Delphi adopt quadratic approximations to approximate activation functions Boemer et al. (2019); Mishra et al. (2020). These methods were proposed to handle only simple model, as non-trivial errors generated by approximation result render the inference results useless in more complex neural networks. On the other hand, the use of high precision approximation guarantees inference accuracy at the cost of higher computational overhead. A common method for precise approximation is Minimax approximation Lee et al. (2022a; 2021b), which combines many small-degree polynomials to approximate non-polynomial functions. Although precise approximation can attain the same accuracy as the backbone model, it leads to much higher costs. Table 1 gives the time cost of the backbone model, their polynomial approximation at various precision, and the corresponding accuracy using state-of-the-art techniques. It can be seen that the PANN on ResNet-20 with precision $2^{-12}$ takes 71.2s to perform inference on the entire CIFAR-10 test set, whereas precision $2^{-8}$ only needs 28.2s. However, the inference accuracy is also significantly lower.
@@
We make two additional remarks here. Firstly, while the overhead of using PANN is high, representing the non-polynomial functions using basic arithmetic or boolean operations is even more expensive. Secondly, we consider PANN as a standalone object, and compare their performance directly. When they are used in combination with advance cryptosystems, the difference will be further enlarged [^1] For instance, in MPC, the function to be computed is first converted into a boolean or arithmetic circuit. If the circuit consists of $n$ gate, the resulting cost of the MPC is $O(n)$ but not necessarily linear in $n$..
@@
Some efforts have been made to increase PANNs’ accuracy while reducing overheads by customizing the model for PANN. For example, CryptoNets and Faster Cryptonets directly train models with low-degree polynomials approximation Gilad-Bachrach et al. (2016); Chou et al. (2018). AutoFHE fine-tunes the model on precise PANN to achieve better performance Ao & Boddeti (2023). However, there are several subtleties involved. Firstly, the polynomials’ derivative is unbounded, making it hard to run gradients descent. Secondly, these models are not generic. Different approximation methods or even different precision for the same method can use different polynomials. This implies that models must be trained separately for each method and precision. Thirdly, in many cases pre-trained model are available, and thus training (or fine-tuning) over PANN can result in much higher training cost and may not be desirable for real-world deployment.
@@
Table 1: Time cost of backbone models (bb) and PANN with different precisions. Numbers in brackets indicate inference accuracy.
@@
To sum up, past researches appear to indicate that a trade-off seems inevitable: Low precision yields high efficiency but low accuracy. High precision yields high accuracy but unaffordable computation costs. The effect of approximation is relatively less understood. This serves as the motivation of this paper. Looking ahead, to better address this problem, as a first step, we separate PANN from the underlying cryptosystem, and investigate how approximation error affects inference accuracy in PANN. Based on this new understanding, we design a more general and efficient methodology to train backbone models so that the corresponding PANN gives reasonable inference accuracy even with approximation at low precision. Furthermore, we highlight that backbone models are still trained with non-polynomial components before they are used to produce PANN. The backbone models trained based on our method do not degrade its inference accuracy.
@@ 1.1 Overview of Our Results
We initiate the investigation on how to reduce the impact on inference accuracy due to approximation errors in PANN. For the ease of writing, we say a neural network $\mathbb{F}$ is sturdy if the PANN of $\mathbb{F}$ has good resistance against approximation errors. In other words, we can use less precise approximation for a sturdy neural network to maintain the inference accuracy. Our investigations lead to the following two interesting findings: (1) The concept of ”sturdiness” is similar to the notion of adversarial robustness (introduced by Projected Gradient Descent Attack (PGD) Kurakin et al. (2016)). Informally speaking, adversarial robustness aims to ensure the neural network is robust to perturbations of the inputs by an adversarial intent on misleading the classifier during the inference stage. We identified the similarity (and difference) between “sturdiness” and “robustness”. In particular, adversarial robustness is related to input perturbation and “sturdiness” is related to perturbation introduced by approximation errors. There are two differences: in adversarial robustness, perturbations are introduced intentionally at input, while for “stuidiness”, perturbations are introduced at every layer with approximation functions by approximation errors, which we referred to as intra-model perturbations. (2) Weight regularization commonly used during the training phase is extremely harmful to model “sturdiness”. Our empirical study shows the following counter-intuitive example: there exists a neural network $\mathbb{F}_{1}$ whose inference accuracy is higher than that of $\mathbb{F}_{2}$, yet the PANN of $\mathbb{F}_{2}$ performs much better than $\mathbb{F}_{1}$ at the same approximation precision. The only difference between $\mathbb{F}_{1}$ and $\mathbb{F}_{2}$ is the use of weight regularization during training of $\mathbb{F}_{1}$.
Based on our findings, we present two solutions to improve “sturdiness” of neural networks. Our first solution is based on the insight that “sturdiness” and robustness are similar in nature. In more detail, our first solution makes use of the technique known as adversarial training (AT), which increases robustness by introducing perturbation during training. However, naively applying perturbations at every layer requiring approximation does not work since it significantly damage the quality of the model. We address this issue by carefully crafting the kind of perturbation to be introduced during training. More concretely, due to the characteristics of the ReLU function, introducing perturbations on the negative part of the inputs during training will barely damage the accuracy of the backbone model. The reason is that information in these negative inputs should be filtered by ReLU function thus they are unimportant to the inference results. Yet, in PANN, these perturbations mimic the “intra-model” perturbations due to the approximation error. Since the perturbations are introduced at the negative part during training and usually corresponds to background information of inputs, we called them as perturbations on irrelevant information in the input background. And we called our method AT-like since it is inspired by adversarial training. Our second solution is based on the findings that weight regularization is harmful to “sturdiness”. Intuitively, we can improve “sturdiness” by removing weight regularization during training. However, weight regularization is crucial in providing generalization and training without it will results in a model with low accuracy. Our second solution is considered a trade-off: we use minimal weight regularization and combine it with Mixup to maintain an acceptable accuracy in the backbone model yet greatly increases its “sturdiness”.
Our two solutions are orthogonal and can be used together during training. We conduct extensive experiments to illustrate the effectiveness of our solutions. In particular, we showed that the resulting PANNs can achieve similar accuracy as state-of-the-art PANNs at a much lower approximation precision.
Our contributions are summarised below:
  • We initiate the study on PANNs and identified similarities and differences between the notion of adversarial robustness and “sturdiness”.
  • We observe that weight regularization can significantly reduce the “sturdiness” of neural networks, thus reducing the corresponding PANN’s accuracy.
  • We propose two solutions to enhance “sturdiness”, namely, an AT-like training method and reduce the use of weight regularization. Combined use of our solutions can significantly enhance “sturdiness”. Our solution is proven effective on various models and datasets. For instance, we achieve the state-of-the-art accuracy on CIFAR-10 with only 40% to 60% time cost (as shown in Figure 1).
@@ 1.2 Outline
The rest of the paper is organized as follows. Section 2 describes the preliminary of our study. In Section 3, we conduct an in-depth analysis on approximation errors in PANN. Section 4 provides solutions to improve low-precision PANN’s accuracy. Section 5 gives the experiment results. Finally, in Section 6, we conclude this paper and discuss potential avenues for future research.

### v2

@@ 1 Introduction
Machine learning (ML) is revolutionising different industries in recent years. Despite its widespread adoption, a key challenge in deploying ML is ensuring data privacy. One promising direction is to employ advance cryptosystems such as multiparty computation (MPC) [1, 2, 3] and homomorphic encryption (HE) [4, 5] during model training and model inference, resulting in privacy-preserving training and privacy-preserving inference. In this paper, we use the term Privacy-preserving machine learning (PPML) to cover both cases.
While in principle these advance cryptosystems can be used to protect sensitive data used in the computation of any functionality, in practice they are designed to support basic arithmetics operations (i.e., addition and multiplication) or basic Boolean operations (i.e., AND and OR) only. This limitation makes it expensive to compute non-polynomial functions such as ReLU, sigmoid, and maxpool, commonly found in ML, and in particular, neural networks. Existing PPML address this limitation by replacing these functions with their polynomial approximations [2, 5, 6, 3]. In this paper, we refer to this type of neural network in which non-polynomial functions have been replaced with polynomials as polynomial approximation of neural network (PANN). Similar to existing MPC and HE based PPML, this paper focuses on privacy-preserving inference since privacy-preserving training using HE and MPC are still considered too inefficient in practice.
Naturally, the (implicit) use of PANN in privacy-preserving inference involves two design considerations, namely, the precision of the polynomial approximation and the accuracy of the PANN (compared with the underlying backbone model). Intuitively, the accuracy of inference cannot be guaranteed if the approximation introduces a large error. For example, early schemes with simple approximation and large errors can handle only simple model and tasks, as non-trivial errors generated by approximation result render the inference results useless in more complex neural networks [4, 7, 8, 1].
On the other hand, the use of high precision approximation guarantees inference accuracy at the cost of higher computational overhead [9, 10]. Although precise approximation can attain the same accuracy as the backbone model, it leads to much higher costs. Table I gives the time cost of PANN and their backbone models, and the corresponding accuracy using state-of-the-art techniques (without fine-tuning on PANN). It can be seen that the PANN on ResNet-20 with error bound $2^{-12}$ takes 71.2s to perform inference on the entire CIFAR-10 test set, whereas that with error bounds $2^{-8}$ only needs 28.2s. However, the inference accuracy is also significantly lower.
@@
We make two additional remarks here. Firstly, while the overhead of using PANN is high, representing the non-polynomial functions using basic arithmetic or boolean operations is even more expensive. Secondly, we consider PANN as a standalone object, and compare their performance directly. When they are used in combination with advance cryptosystems, the difference will be further enlarged [^1] For instance, in MPC, the function to be computed is first converted into a boolean or arithmetic circuit. If the circuit consists of $n$ gates, the resulting cost of the MPC is $O(n)$ but not necessarily linear in $n$..
Some efforts have been made to increase PANNs’ accuracy while reducing overheads by customizing the model structure for PANN or fine-tuning (training) the model on PANN [4, 2, 11, 3] and achieve satisfactory performance.
@@
However, there are several subtleties involved. Firstly, the polynomials’ derivative is unbounded, making it hard to run gradients descent. Secondly, these models are not generic. Different approximation methods or even different precision for the same method can use different polynomials. This implies that these models must be trained separately for each method and precision. Thirdly, in many cases, popular model structures have been widely studied and provide pre-trained models, and thus training (or fine-tuning) over PANN or non-generic model structures can result in much higher validation cost and may not be desirable for real-world deployment.
@@
TABLE I: Time cost of backbone models (bb) and PANN with different error bound. Numbers in brackets indicate inference accuracy.
@@
To sum up, past researches appear to indicate that a trade-off seems inevitable: Low precision yields high efficiency but low accuracy. High precision yields high accuracy but unaffordable computation costs. The effect of approximation is relatively less understood. This serves as the motivation of this paper. Looking ahead, to better address this problem, as a first step, we separate PANN from the underlying cryptosystem, and investigate how approximation error affects inference accuracy in PANN.
@@ 1.1 Overview of Our Results
We initiate the investigation on how to reduce the impact on inference accuracy due to approximation errors in PANN. For the ease of writing, we say a neural network $\mathbb{F}$ is sturdy if the PANN of $\mathbb{F}$ has good resistance against approximation errors. In other words, we can use less precise approximation for a sturdy neural network to maintain the inference accuracy.
Specifically, we bound the lower bound of increased loss resulting from approximation errors and obtain the following two interesting results:
  • Approximation errors on negative inputs of ReLU lead to larger loss than those on positive inputs of ReLU;
  • The increased loss resulting from approximation errors will increase with larger weight decay and more training epochs.
To substantiate our analysis, we conduct tests and have observed the results across various model structures, datasets, and approximation methods.
Based on our findings, we present two solutions to improve “sturdiness” of neural networks, namely, the negative inputs leakage method and reducing the use of weight regularization. Our first solution is based on the insight that approximation errors on ReLU’s negative inputs lead to more loss. In more detail, our first solution introduces perturbations on the negative inputs of ReLU during baseline training. Our second solution is based on the findings that weight regularization is harmful to “sturdiness”. Intuitively, we can improve “sturdiness” by removing weight regularization during training. However, weight regularization is crucial in providing generalization. Our second solution is considered a trade-off: we use minimal weight regularization and combine it with Mixup to maintain an acceptable accuracy in the backbone model yet greatly increase its “sturdiness”. We conduct extensive experiments to illustrate the effectiveness of our solutions. For instance, we achieve the state-of-the-art non-fine-tuning accuracy on CIFAR-10 with only 40% to 60% time cost (as shown in Figure 1). We improve the accuracy of the state-of-the-art ReLU replacement scheme [2] by 3%. Compared to previous schemes, our methods have the following advantages:
  • Our obtained models are based on generic model structures rather than being specifically fine-tuned for particular PANNs. That means the same model can perform as the backbone (or pre-trained) model for various approximation precision and methods to enhance their performance without adjustment. This lowers the training cost and ensures better compatibility with other approaches.
  • Our training methods are also general without restricting to particular model structures or approximation methods. This allows for a variety of fine-tuning-based schemes to be further optimized with our approaches.
@@ 1.2 Outline
The rest of the paper is organized as follows. Section 2 gives the theorem and explanation of how approximation errors affect PANN. Section 3 introduces the related works. Section 4 describes the preliminary of our study. Section 5 provides solutions to improve low-precision PANN’s accuracy. Section 6 gives the experiment results. Finally, in Section 7, we conclude this paper and discuss potential avenues for future research.
(a) ResNet-20
(b) Shufflenetv2
(c) DLA-34
(d) Mobilenetv2
Figure 1: Testing loss increment (compared to baseline) of PANN [5] approximating ReLU with positive inputs (pos) and negative inputs (neg) on models trained with weight decay (wd) 0 and 1e-4. It shows that weight decay can amplify the loss increment caused by approximation. Additionally, approximating ReLU with negative inputs results in more loss compared to approximating ReLU with positive inputs.
2 Effects of Approximation Errors on PANN
In this section, we present our main theorem and discuss the factors that affect the loss increment caused by approximation errors. Since for most PANN schemes, activation functions, especially the ReLU function, are the main components requiring approximation, we will focus specifically on the approximation of ReLU functions.
Considering a two-layer neural network with ReLU activation functions. The input point is a $d$-dimentional vector $\textbf{{x}}\in\mathbb{R}^{d}$. The weights is $\textbf{{W}}\in\mathbb{R}^{n*d}$, let $\textbf{{z}}=\textbf{{W}}\textbf{{x}}$ and $\textbf{{y}}=\operatorname{ReLU}(\textbf{{z}})$. For a loss function $L$, weight decay parameter $\lambda$, and learning rate $\eta$, we define the loss function for SGD with with L2 regularization at epoch $t$ as:
$f(\textbf{{W}}_{t})=L(\textbf{{W}}_{t})+\frac{\lambda}{2\eta}\|W_{t}\|^{2}$ (1)
For the ease of writing, we use $h(\operatorname{ReLU}(z))=f(\textbf{{W}},\textbf{{X}})$ to represent the same loss but with inputs $\operatorname{ReLU}(z)$ where $z$ is a value in z, such that the loss with error $\epsilon$ is $h(\operatorname{ReLU}(z)+\epsilon)$. We use $\Delta h$ to represent the increased loss caused by error $\epsilon$ (Note that $\Delta h=\Delta f$). Now we give two theories and the corresponding assumptions.
Approximation errors on the negative inputs of ReLU increase more loss than those on positive inputs
A question that remains unexplored is which part of approximation errors lead to more accuracy reduction (or a higher loss). Here we give Theorem 2.1 which indicates that errors on negative inputs of ReLU result in larger loss and therefore poorer accuracy.
Theorem 2.1.
Let two convex function $h_{1}:\mathbb{R}\to\mathbb{R}$ and $h_{2}:\mathbb{R}\to\mathbb{R}$ differentiable in $(-\infty,0)\cup(0,\infty)$. Let $\overline{z}_{1}\in\mathbb{R}$ and $\overline{z}_{2}\in\mathbb{R}$ and $\overline{z}_{1}<0,\overline{z}_{2}>0$, such that $h_{1}^{\prime}(\overline{z}_{1}))=h_{2}^{\prime}(\overline{z}_{2}))=0$. Let $\epsilon\in\mathbb{R}$ be a small value, we have:
$\lim_{\epsilon\to 0}(\Delta h_{1}-\Delta h_{2})=\epsilon\cdot\mathop{\operatorname{inf}}_{z>0}\phi_{0,h_{1}}(z)$ (2)
The theorem suggests that when approximating ReLUs with negative inputs, the loss is at least as large as approximating those with positive inputs. This implies that it’s beneficial to reduce the approximation of ReLU with negative inputs or to develop specific training strategies for negative inputs.
In Figure 1, we present the increase in loss when only approximating ReLU with negative (neg) or positive (pos) inputs. It illustrates the phenomenon that approximating ReLUs with negative inputs results in more loss than approximating ReLUs with positive inputs (when errors are small). This can be observed in various model structures and provides evidence for our theorem.
Weight decay enlarges the loss increasing resulting from approximation error as training progresses
Another important question is which factors in backbone model training affect the “sturdiness”. Now we provide Theorem 2.2, which proves that weight decay is responsible for weak sturdiness.
Theorem 2.2.
Let function $L:\mathbb{R}\to\mathbb{R}$ be a $\mathcal{L}$-smooth function and $h$ be a convex function with $h^{\prime}(\operatorname{ReLU}(\overline{z}))=0$. Assume $h$ is differentiable in $(-\infty,0)\cup(0,\infty)$. Let $\epsilon\in\mathbb{R}$ be a small value. $\mathbb{E}[\nabla L(\textbf{{W}},X)-\nabla L(\textbf{{W}})]=0$, $\mathbb{E}[\|\nabla L(\textbf{{W}},X)-\nabla L(\textbf{{W}})\|^{2}]\leq\sigma^{2}$, $\|\nabla l(\textbf{{W}})\|\leq G$ for any W, and $\eta\leq\frac{C}{\sqrt{t+1}}$. If the model is trained $t^{\prime}+1$ epochs after it has reached the stationary point $\textbf{{W}}^{*}$ where $L(\textbf{{W}}^{*},X)=L^{\star}$, the increased loss $\Delta L=h(\operatorname{ReLU}(\overline{z}+\epsilon))-h(\operatorname{ReLU}(\overline{z}))$ caused by error $\epsilon$ is bounded by:
$||\epsilon||^{2}\cdot\mathbb{E}[||\mathop{\operatorname{inf}}_{z>\overline{z}}\phi_{\overline{z}}(z)||^{2}]\leq\mathbb{E}[||\Delta h||^{2}]$ (3)
where
$\operatorname{sup}\mathbb{E}[||\operatorname{inf}_{z>\overline{z}}\phi_{\overline{z}}(z)||^{2}]\leq||h^{\prime}_{+,t}(\overline{z})||^{2}+\frac{t^{\prime}(C_{1}+C_{2})}{d\sqrt{t+t^{\prime}+1}}$
and:
$C_{1}$ $=\frac{L(\textbf{{W}}_{0})-L^{\star}}{C}$
$C_{2}$ $=C(\mathcal{L}+\lambda)((G+\sup(\lambda\|W\|))^{2}+\sigma^{2}),$ (4)
The theorem demonstrates that the loss will increase with larger weight decay $\lambda$ and epoch $t^{\prime}$. We have observed and confirmed the effect of these two factors on PANN.
An example of PANN [5] on models trained with different weight decay is shown in Figure 2. It illustrates that though the backbone (bb) models have similar accuracy, the accuracy of PANN reduces when weight regularization ($\lambda$) is applied as the epochs ($t^{\prime}$) increase. Larger weight decay can make this reduction quicker ($\lambda$ increases). Furthermore, Figure 1 illustrates the increase of testing loss of PANN [5] with different backbone and different weight decay, indicating that weight decay can elevate the testing loss in all cases. (Still, note that the error is small.)
Figure 2: Backbone models (bb) trained with different weight decay (wd) have similar accuracy (dot line). However, large weight decay can reduce their accuracy on PANN as the epoch increases. (ResNet-18, CIFAR-100, precision $2^{-9}$)
This phenomenon can be observed in different model structures, datasets, approximation error bounds (Table II, Table III), and approximation methods. We believe the above theorems and analysis provide a plausible explanation on how approximation errors reduce the accuracy of PANN. (Detailed proof of the two theorems is put in Appendix B.)
An intuitive explanation
Here we give an intuitive explanation based on the definition of stationary point to help understand the two theorems. A stationary point is where the function’s derivative is zero, and is a necessary condition for minimizing training losses. For stochastic gradient descent (SGD) method without weight regularization, we have $f=L$. The model weights for epoch $t+1$ are updated by:
$\textbf{{W}}_{t+1}=\textbf{{W}}_{t}-\eta\frac{\partial L(\textbf{{W}}_{t})}{\partial\textbf{{W}}_{t}},$ (5)
then model training is searching for the smallest loss which have $\frac{\partial L(\textbf{{W}}_{t})}{\partial\textbf{{W}}_{t}}=\mathbf{0}$.
Let’s focus on a single input sample $x_{1}$ and the $j$-th column of W (represented as $\textbf{{W}}_{j}$). Let $z$ be a single value in vector z and $y$ be the corresponding value in y that satisfies $y=\operatorname{ReLU}(z)=\operatorname{ReLU}(\textbf{{W}}_{j}x_{1})$. Assume a non-zero stationary point $\textbf{{W}}^{*}$ is found at epoch $t$. For $\lambda=0$, by combining Equation 1, we can get:
$\frac{\partial L}{\partial\textbf{{W}}^{*}_{j}}=\frac{\partial L}{\partial y}\frac{\partial y}{\partial z}\frac{\partial z}{\partial\textbf{{W}}^{*}_{j}}=0$ (6)
In this condition, the neural network has minimal loss and is the most stable because small input changes will have little effect on the outputs. However, for PANN with approximation errors, the value $y$ is changed to:
$\widetilde{y}=y+\epsilon_{relu}(z)$ (7)
This leads to a difference between PANN and the backbone (or pre-trained model) that a stationary point in the original model may not be the stationary point in PANN. We discuss it in two cases, ReLU with negative or zero inputs ($z<0$) and ReLU with positive inputs ($z>0$).
For $z<0$, we have $\frac{\partial y_{-}}{\partial z}=\frac{\operatorname{ReLU}(z)}{\partial z}=0$. Due to the property of ReLU, $L$ is non-differential at the point $z=0$. So we have $\frac{\partial y_{-}}{\partial z}\neq\frac{\partial y_{+}}{\partial z}$.
From Equation 6 we can know $\frac{\partial L}{\partial\textbf{{W}}^{*}_{j}}=0$ and the stationary point holds for all $z\leq 0$ in the backbone model. However, in PANN $\operatorname{ReLU}(z)$ is replaced with $\operatorname{ReLU}(z)+\epsilon$. Assume $\frac{\partial L}{\partial y}\frac{\partial z}{\textbf{{W}}^{*}_{j}}\neq 0$ (which is the most cases), then $\textbf{{W}}^{*}$ may not be a stationary point of PANN, because from 7 we can get:
$\frac{\partial\widetilde{y}}{\partial z}=\frac{\partial y_{+}}{\partial z}+\frac{\partial\epsilon}{\partial z}$ (8)
By combining Equation 6 and 8, we get:
$\frac{\partial\widetilde{L}}{\textbf{{W}}^{*}_{j}}=\frac{\partial L}{\partial y}\frac{\partial\epsilon}{\partial z}\frac{\partial z}{\textbf{{W}}^{*}_{j}}$ (9)
As long as $\frac{\partial y_{+}}{\partial z}+\frac{\partial\epsilon}{\partial z}\neq 0$, we have $\frac{\partial\widetilde{L}}{\textbf{{W}}^{*}_{j}}\neq 0$, which means $W^{*}$ is not the stationary point of PANN. The loss is not a minimum and is not sufficiently stable.
Differently, for $z>0$, we have $\frac{\partial y}{\partial z}=1$ in backbone model. By combining Equation 6 we have:
$\frac{\partial L}{\partial y}\frac{\partial z}{\textbf{{W}}^{*}_{j}}=0$ (10)
For PANN, from Equation 7 we can get:
$\frac{\partial\widetilde{y}}{\partial z}=\frac{\partial y}{\partial z}+\frac{\partial\epsilon}{\partial z}=1+\frac{\partial\epsilon}{\partial z}$ (11)
Because ReLU is differentiable for positive inputs, when the approximation error is small enough, the neural network can be seen as linear and $\frac{\partial L}{\partial y}\frac{\partial z}{\textbf{{W}}^{*}_{j}}$ remains unchanged. By combining Equation 6 and 11, we have:
$\frac{\partial\widetilde{L}}{\partial W_{j}}=(1+\frac{\partial\epsilon}{\partial z})\frac{\partial L}{\partial y}\frac{\partial y}{\partial z}\frac{\partial z}{\textbf{{W}}^{*}_{j}}=0$ (12)
That means the backbone model and PANN share the same stationary point when errors are added to positive inputs ($z_{j}>0$) of ReLU, which reduces the effect of approximation errors. Besides, even if the weight was driven away from $W^{*}$ during the training, an item $-\frac{\partial L(W,x_{1})}{\partial W}$ in weight updating will fix the stationary point towards $W^{*}$.
When weight decay is involved, it has been proved that weight decay can lead to large gradient norms [12]. It may lead to the following problems: (1) Increase the norm of $z$, thus increasing the amplitude of some noise proportional to the $||z||$; (2) Increase $\frac{\partial L}{\partial z_{+}}$ when $z<0$. Because the stationary points $\frac{\partial L}{\partial z_{-}}=0$ always hold, $\frac{\partial L}{\partial z_{+}}$ may be trained towards away from the stationary point. Weight decay will accelerate this process (by destroying patterns established before $z$ is negative).
3 Related Works and Background
3.1 Approximation Methods for Neural Networks
Replacing non-polynomial functions with polynomials has become a desirable tool for applying cryptography to neural networks. Early studies often adopt simple (low-degree) polynomials in approximation. For example, CryptoNets and HCNN utilized the square function to replace activation functions [4, 7]. CryptoDL used degree 2 and degree 3 polynomials to approximate the ReLU function [13], while Faster Cryptonets leveraged minimax approximation with degree two [14]. nGraph-HE and Delphi adopt quadratic approximations to approximate activation functions [8, 1]. These methods can handle only simple model (i.e. less than 10 layers) and tasks (i.e. classification on MINIST) because the non-negligible errors accumulated throughout layers can seriously affect the model outputs.
To address this, further studies proposed highly precise approximation to reduce the approximation error and can achieve the same accuracy as the backbone model [5, 15]. A common method for these precise approximations is Minimax approximation [9, 10, 16], which combines many small-degree polynomials for approximation to reduce the computation cost. However, the reduced computational overhead of precise approximation is still unbearable. Increasing PANNs’ accuracy while reducing overheads remains an important issue.
3.2 Fine-tuning-based Solutions
When seeking to improve the performance of PANN, many studies fine-tune (or train) models directly on PANN. Earlier studies such as CryptoNets and Faster Cryptonets directly train models on neural networks with low-degree polynomials activation functions [4, 14]. The models obtained are essentially new models with a customized design for PANN (both in terms of structure and parameters) rather than being approximations of commonly used models. The limitations of simple polynomials activation and the customized structure make these models hard to tackle complex tasks. To better suit practical applications, recent studies tend to input pre-trained models as a starting point and then fine-tune them on PANN. For example, SNL and AutoReP fine-tune models after replacing a partial of ReLU functions with polynomials with one or two degree [17, 2]. AutoFHE fine-tunes models on precise PANN to achieve better performance and lighten the requirements for precision [11]. Another study fine-tuning models after carefully adjusting the model structure for low-degree polynomial approximations has also obtained a high accuracy [3].
Despite its efficiency in improving PANN accuracy, several issues are involved. Firstly, polynomials are not appropriate for training when used as activation functions. The derivative of these polynomials is unbounded, which can lead to unstable values during gradient descent and get unexpected results (i.e. the parameter instability issues in AutoReP [2]). Secondly, a fine-tuned model is customized for the weights and polynomials it’s trained on. The model must be retrained whenever you want to adjust the PANN’s approximation precision and overhead (discussed in Section 4.2), which will result in additional training costs. Thirdly, generic model structures and activation functions are usually well-validated. Designing extra network structures for PANN results in higher design and test costs and is undesirable for other applications. Therefore, our study prefers to solve PANN’s accuracy reduction problems on generic model structures crossing different approximation methods and precision requirements without fine-tuning.
3.3 ReLU Replacement Schemes
Reducing the number of activation functions in PANN can significantly reduce computation costs since they are the main part requiring approximation and cryptographic operations. SNL first achieves this by designing a parametric ReLU (PReLU) and applying a loss function to gradually adjust ReLU functions to linear functions [17]. This method can reduce the ReLU number in a ResNet-18 model from 491.52K to 49.9K while maintaining 73.75% accuracy (baseline 77.8% on CIFAR-100). AutoRep improves the method by introducing parameterized discrete indicator function, hysteresis loop update function, and distribution-aware polynomial approximation [2], which improves the accuracy to 75.48%.
Though ReLU replacement can significantly reduce the computational overhead of PANN, two factors reduce the accuracy of obtained PANNs in this process. First, the structure of the pre-trained model is different from PANN. However, the pre-training parameters are not resistant to perturbations caused by structural changes. Although this effect can by reduced by fine-tuning, there is still much room for improvement. Second, the activation functions during backpropagation and the activation functions in the final obtained models are also different. This further reduces PANN’s accuracy.
Actually, the two factors point to the same problem that the model is not resistant to approximation errors (we say “poor sturdiness”). While our solutions, which increase model sturdiness based on general structures, can serve as a supplementary to ReLU replacement schemes. By taking our models as pre-trained and teacher models, the performance of SNL and AutoRep with different ReLU counts can all be further improved.
3.4 ReLU Protocol based on Truncation
Bicoptor [18] first proposed an efficient three-party computation (3PC) protocol for ReLU based on truncation and Bicopter2.0 [19] further improved it. For a fix-point input $x$ with $l_{x}$ bit length, the protocol will right-shift the complement code of $x$ for $l_{x}$ times with truncation and get $l_{x}$ shares. By securely checking if zero appears in these shares, the protocol will determine the sign of $x$ (i.e., $x=0b00001010$ right shift 4 bit will get 00) and enable the computing of $\operatorname{ReLU}(x)$. Since this process involves transferring $lx$ truncation results, the communication cost will increase with $l_{x}$ ($(+2)l$ bit for Bicoptor, $(l_{x}+1)(l_{x}+1)$ for Bicoptor2.0).
This means choosing a suitable (small) $l_{x}$ in ReLu protocol can effectively reduce the communication overhead at the cost of computation precision. We considered this error also a kind of approximation error. Though it’s different from polynomial approximation, the effects is still in line with the results of our analyses and can apply our solutions.
3.5 Deep Learning with Differential Privacy
Differential privacy (DP) protects the privacy of individual tuples in a database [20, 21, 22, 23]. It ensures that attackers cannot infer the membership of any tuple from the released data while the statistical properties are preserved for usage. By applying DP in machine learning, models can be trained without exposing sensitive information [24, 25, 26, 27]. DP-based machine learning does not require time-consuming cryptographic computation and thus is highly practical. However, several serious issues are involved. Firstly, the statistics information of the dataset is also valuable. Especially when the data collection is costly, data owners may be reluctant to disclose any information. Secondly, DP cannot absolutely protect user privacy. Attackers can still get information about some users, though with a low probability. Thirdly, DP aims to hide data in a database, which should contain a large amount of data. This requirement cannot be satisfied in scenarios like privacy-preserving inference.
While cryptographic methods can protect the information of all inputs (both statistical information and individual tuples) and have no requirements for the dataset. In more detail, cryptographic schemes can ensure that every ciphertext in the computation process does not reveal the plaintext information if the attacker doesn’t have the key. Cryptographic techniques such as FHE and MPC can compute the data in the ciphertext format, and the result is generated and stored in ciphertext, which can only be revealed after decryption.
3.6 Side-effects of Weight Decay
Weight decay is a powerful regularization technique that has been very widely used in training deep neural networks. It can improve model generalization and accuracy [28]. However, recent studies have shown that this benefit comes with side effects [29, 12]. Weight decay has been proved to increase bad minima when applied in training neural networks [29]. Further study has proved that weight decay can lead to large gradient norms at the final phase of training, which often indicates bad convergence and poor generalization [12]. This large gradient norm problem is significant in the presence of scheduled or adaptive learning rates and has been recognized as harmful to neural networks. In our study, we have also confirmed that weight decay reduces the sturdiness of the model, which to some extent, supports these arguments.